# Understanding Code Understandability Improvements in Code Reviews

### v1

@@ 6.2 Estimating and classifying code understandability
Fakhoury et al. [27] investigated whether state-of-the-art understandability models are able to capture understandability (readability in their paper) improvements as explicitly tagged by open source developers in commit messages. The authors analyzed 548 commits from 63 engineered Java projects that explicitly mention understandability improvements in the commit message. They then used three state-of-the-art understandability models to measure how much these commits affected code understandability. They found that the models did not capture understandability improvements in most cases. Additionally, they used static analysis tools to investigate what type of understandability changes were applied in the commits related and unrelated to code understandability. They discovered that the former frequently fix concerns related to imports, white spaces, and braces, whereas the latter tend to introduce them.
@@
Roy et al. [28] proposed a model to detect code understandability (readability in their paper) improvements made by developers in real-world scenarios using commit histories from open-source software projects by extending the dataset of Fakhoury et al. [27]. They also included commits unrelated to code understandability improvements to create an oracle. The authors employed seven source code analysis tools to collect metrics before and after each commit at the file level. By comparing these metrics, they identified key differences that served as the features for their model. The model achieved a precision of 79.2% and a recall of 67% on the test set.
@@
Buse and Weimer [21] proposed a method for measuring code understandability (code readability in their paper) based on human notions of understandability. They collected feedback from 120 human annotators on their understanding of 100 code snippets, correlating this feedback with source code features. This data was then used to train a machine learning algorithm to classify code understandability. The resulting model predicted understandability judgments with an 80% success rate. Posnett et al. [80] proposed a simplification of Buse and Weimer’s [21] model. They leverage source code size metrics and Halstead metrics [81]. Using the dataset of Buse and Weimer [21] to train and evaluate their model, they obtained better results using just three features in small code snippets: volume, entropy and number of lines.
@@
Scalabrino et al. [15] analysed the correlation between 121 metrics related to code, documentation, and developers and six different proxies of code understandability, based on 444 human evaluations. The study revealed that none of the individual metrics strongly correlated with the proxies for code understandability. However, combining multiple state-of-the-art metrics resulted in models that showed a slight improvement in the regression performance of two proxies, Timed Actual Understandability (Correlation: +0.07; Mean absolute error: -0.02) and Actual Understandability (Correlation: +0.02; Mean absolute error: +0.00).
@@
Some other studies [83, 84, 85] have proposed techniques for automating assessment through advanced methods, including convolutional neural networks. Mi et al. [83] introduced IncepCRM, an model that autonomously extracts multi-scale features from code, aided by human annotations to enhance accuracy. In a distinct approach [84], an 83.8% accuracy is achieved, surpassing prior machine-learning models. This is realized by transforming source code into matrices for convolutional neural networks, culminating in DeepCRM, an architecture of three networks trained on diverse preprocessed data. Similarly, in another study [85], a graph-based method is adopted, parsing source code into a graph containing abstract syntax tree (AST), along with control and data flow edges, preserving semantic structural insights. This approach attains 72.5% and 88% accuracy in three-class and two-class classifications, respectively. Our analysis of inline code reviews, focusing on developers’ perspectives regarding code understandability, provides valuable insights for the referenced works. The research conducted by Mi et al. [83] could further benefit by incorporating insights from our study to refine their IncepCRM model. Our findings might aid in identifying specific aspects of code understandability highlighted during reviews, enabling more effective feature extraction for multi-scale representation. Similarly, the strategy proposed by the author in [84] could gain from our insights by considering the nuances of developer concerns about code understandability in the design of novel representation strategies. Integrating our unique dataset and analytical framework could enhance these approaches, contributing to advancements in code understandability evaluation.
@@ 6.3 Understandability and Code Reviews
The work of Pascarella et al. [88] aims to understand the information necessary for reviewers to conduct a proper code review. It suggests ways in which research and tool support can enhance the effectiveness and efficiency of developers in their role as reviewers. The study analyzed 900 code review comments from three large open-source projects and identified seven primary information needs of reviewers, such as understanding the uses of methods and variables declared or modified within the code. In a similar vein, Ebert et al. [60] investigates the causes and consequences of confusion in code reviews and how developers typically handle such issues. The study involved a survey of 54 valid responses and a manual evaluation of 307 code review comments (both inline and general). The authors identified 30 sources of confusion, 14 ways that confusion impacts code review, and 13 coping strategies to deal with confusion in code reviews. The research revealed that developers often experience confusion during code reviews when dealing with long or complex code changes and changes that address multiple issues.
@@
Other studies have explored approaches to automate code review, in the generation of patches [89, 90] and of code review comments [91]. The former approach leverages LLMs and the latter employs information retrieval. Neither of them attempts to distinguish the purposes of the code review comments. Likewise, the study by Fregnan et al. [92] evaluated what extent a machine learning-based technique can automatically classify review changes. The researchers classified 1,504 review changes manually and then evaluated three different machine learning algorithms to see if they could classify review changes automatically at two different levels of detail. Their findings indicate that machine learning can be effectively used to classify review-induced changes. This paper is complementary to these approaches in that it narrows down the scope of investigation to code understandability improvements. It shows relevance of code review comments aiming to improve understandability, highlights the high level problems that these comments address, shows how many of the corresponding improvements materialize, and indicates that they are often accepted and rarely reversed. It also provides evidence that automating the application of improvements involves a variety of issues, targets different languages, natural and artificial, and in some cases can be solved easily whereas in others requires a deeper understanding.
@@ Acknowledgements
This work was partially supported by the the Swedish Foundation for Strategic Research (SSF) and the Wallenberg Artificial Intelligence, Autonomous Systems and Software Program (WASP) funded by Knut and Alice Wallenberg Foundation; and INES (www.ines.org.br), CNPq grant 465614/2014-0, FACEPE grants APQ-0399-1.03/17 and APQ/0388-1.03/14, CAPES grant 88887.136410/2017-00.

### v2

@@ 6.2 Estimating and classifying code understandability
Fakhoury et al. [27] investigated whether state-of-the-art understandability models are able to capture understandability (readability in their paper) improvements as explicitly tagged by open source developers in commit messages. The authors analyzed 548 commits from 63 engineered Java projects that explicitly mention understandability improvements in the commit message. They then used three state-of-the-art understandability models to measure how much these commits affected code understandability. They found that the models did not capture understandability improvements in most cases. Additionally, they used static analysis tools to investigate what type of understandability changes were applied in the commits related and unrelated to code understandability. They discovered that the former frequently fix concerns related to imports, white spaces, and braces, whereas the latter tends to introduce them.
@@
Roy et al. [28] proposed a model to detect code understandability (readability in their paper) improvements made by developers in real-world scenarios using commit histories from open-source software projects by extending the dataset of Fakhoury et al. [27]. They also included commits unrelated to code understandability improvements to create an oracle. The authors employed seven source code analysis tools to collect metrics before and after each commit at the file level. By comparing these metrics, they identified key differences that served as features for their model. The model achieved a precision of 79.2% and a recall of 67% on the test set.
@@
Buse and Weimer [21] proposed a method for measuring code understandability (code readability in their paper) based on human notions of understandability. They collected feedback from 120 human annotators on their understanding of 100 code snippets, correlating this feedback with source code features. This data was then used to train a machine learning algorithm to classify code understandability. The resulting model predicted understandability judgments with an 80% success rate. Posnett et al. [80] proposed a simplification of Buse and Weimer’s [21] model. They leverage source code size metrics and Halstead metrics [81]. Using the dataset of Buse and Weimer [21] to train and evaluate their model, they obtained better results using just three features in small code snippets: volume, entropy, and number of lines.
@@
Scalabrino et al. [15] analyzed the correlation between 121 metrics related to code, documentation, and developers and six different proxies of code understandability, based on 444 human evaluations. The study revealed that none of the individual metrics strongly correlated with the proxies for code understandability. However, combining multiple state-of-the-art metrics resulted in models that showed a slight improvement in the regression performance of two proxies, Timed Actual Understandability (Correlation: +0.07; Mean absolute error: -0.02) and Actual Understandability (Correlation: +0.02; Mean absolute error: +0.00).
@@
Some other studies [83, 84, 85] have proposed techniques for automating assessment through advanced methods, including convolutional neural networks. Mi et al. [83] introduced IncepCRM, a model that autonomously extracts multi-scale features from code, aided by human annotations to enhance accuracy. In a distinct approach [84], an 83.8% accuracy is achieved, surpassing prior machine-learning models. This is realized by transforming source code into matrices for convolutional neural networks, culminating in DeepCRM, an architecture of three networks trained on diverse preprocessed data. Similarly, in another study [85], a graph-based method is adopted, parsing source code into a graph containing abstract syntax tree (AST), along with control and data flow edges, preserving semantic structural insights. This approach attains 72.5% and 88% accuracy in three-class and two-class classifications, respectively. Our analysis of inline code reviews, focusing on developers’ perspectives regarding code understandability, provides valuable insights for the referenced works. The research conducted by Mi et al. [83] could further benefit by incorporating insights from our study to refine their IncepCRM model. Our findings might aid in identifying specific aspects of code understandability highlighted during reviews, enabling more effective feature extraction for multi-scale representation. Similarly, the strategy proposed by the author in [84] could gain from our insights by considering the nuances of developer concerns about code understandability in the design of novel representation strategies. Integrating our unique dataset and analytical framework could enhance these approaches, contributing to advancements in code understandability evaluation.
@@ 6.3 Understandability and Code Reviews
The work of Pascarella et al. [88] aimed to understand the information necessary for reviewers to conduct a proper code review. It suggests ways in which research and tool support can enhance the effectiveness and efficiency of developers in their role as reviewers. The study analyzed 900 code review comments from three large open-source projects and identified seven primary information needs of reviewers, such as understanding the uses of methods and variables declared or modified within the code. In a similar vein, Ebert et al. [60] investigated the causes and consequences of confusion in code reviews and how developers typically handle such issues. The study involved a survey of 54 valid responses and a manual evaluation of 307 code review comments (both inline and general). The authors identified 30 sources of confusion, 14 ways that confusion impacts code review, and 13 coping strategies to deal with confusion in code reviews. The research revealed that developers often experience confusion during code reviews when dealing with long or complex code changes and changes that address multiple issues.
@@
Other studies have explored approaches to automate code review, in the generation of patches [89, 90] and of code review comments [91]. The former approach leverages LLMs and the latter employs information retrieval. Neither of them attempts to distinguish the purposes of the code review comments. Likewise, the study by Fregnan et al. [92] evaluated what extent a machine learning-based technique can automatically classify review changes. The researchers classified 1,504 review changes manually and then evaluated three different machine learning algorithms to see if they could classify review changes automatically at two different levels of detail. Their findings indicate that machine learning can be effectively used to classify review-induced changes. This paper is complementary to these approaches in that it narrows down the scope of investigation to code understandability improvements. It shows the relevance of code review comments aiming to improve understandability, highlights the high-level problems that these comments address, shows how many of the corresponding improvements materialize, and indicates that they are often accepted and rarely reversed. It also provides evidence that automating the application of improvements involves a variety of issues, targets different languages, natural and artificial, and in some cases can be solved easily whereas in others requires a deeper understanding.
@@ Acknowledgements
This work was partially supported by INES (www.ines.org.br), CNPq grant 465614/2014-0, FACEPE grants APQ-0399-1.03/17 and APQ/0388-1.03/14, CAPES grant 88887.136410/2017-00, the Swedish Foundation for Strategic Research (SSF), and the Wallenberg Artificial Intelligence, Autonomous Systems and Software Program (WASP) funded by Knut and Alice Wallenberg Foundation.